OpenStudy (anonymous):

A running back carries the ball five times. Find total total yardage if he carries it 8, -2, -1, 16, & 7 yards.

OpenStudy (anonymous):

8 -2 = 6 6-1= 5 5+16 = 21 21 + 7 = 28